If a student jumps straight up to a height of 0.440m, what is his initial speed as it leaves the ground? How long

is he in the air?


1
Expert's answer
2022-03-02T17:00:08-0500

"v=\sqrt{2gh}=2.9~\frac ms,"

"t=2\sqrt{\frac{2h}g}=0.6~s."